Arthur weighs 34lbs more than Lily. Their combined weight is 180lbs less than four times Lily's weight. How much does Lily weigh?
Answer by stanbon(75887)

Arthur weighs 34lbs more than Lily. Their combined weight is 180lbs less than four times Lily's weight. How much does Lily weigh?
-------
A = L + 34
A + L = 108
----
Substitute for "A" and solve for "L":
(L+34) + L = 108
2L = 74
Lily's weight = 37 lbs.
==================
Cheers,
Stan H.
===================